# Non-universal surface magnetoelectric response in antiferromagnetic topological insulators

### v2

@@ Abstract
In the main text we demonstrated that, due to the symmetry of Eq. (S9), the $\mathcal{N}>0$ Landau level subspaces do not contribute to the electronic polarization. Thus we restrict focus to the Hamiltonian in the $\mathcal{N}=0$ subspace, Eq. (S12), which takes the form of a generalized SSH model. That is, when we consider an infinite number of layers (i.e. implement periodic boundary conditions) in Eq. (S1), the result takes a tight-binding form [15] in the stacking ($\hat{\bm{z}}$) direction. And when this 3D effective model is projected into the $\mathcal{N}=0$ subspace, the result is a one-dimensional (1D) tight-binding model. The ground state polarization in our low-energy 3D model has one bulk component, namely $P^{z}$, which can be calculated using expressions derived in the modern theory of polarization [38, 39]. The $\mathcal{N}=0$ contribution to $P^{z}$ is proportional to the 1D bulk polarization of Eq. (S12) at half-filling, namely
$P_{\text{SSH}}^{z}=e\sum_{n=1}^{2}\int_{\text{BZ}}\frac{dk_{z}}{2\pi}\xi^{z}_{nn}(k_{z}),$ (S16)
@@
where $\xi^{z}_{nm}(k_{z})\equiv i\innerproduct{u_{n,k_{z}}}{\frac{\partial}{\partial k^{z}}u_{m,k_{z}}}$ are band components of the Berry connection induced by a choice of Bloch energy eigenvectors with cell-periodic parts |$\ket{u_{n,k_{z}}}$. Inversion symmetry of the model dictates that $P_{\text{SSH}}^{z}\text{ mod }e\in\{0,e/2\}$ [60]. A set of analytical eigenvectors of Eq. (S12) can be found and the components differentiated with respect to $k_{z}$ in an effort to obtain $\xi^{z}_{nn}$.
@@
One issue remains: how do we take the $k_{z}$-derivative of the basis states with respect to which the Hamiltonian matrix (S12) is written? Typically in phenomenological tight-binding models those basis states are taken to be such that the corresponding orbitals are highly localized in real-space and centered at the ionic positions. Often the derivatives of these basis states do not affect the results of calculations of topological quantities (e.g., polarization of the usual SSH model or the Chern number of the Haldane model) and one can get away with taking the basis states to be independent of $\bm{k}$. However, in the present case accounting for these derivatives becomes important. For if we assume the basis states to be independent of $k_{z}$ then we find, in general, that $P_{\text{SSH}}^{z}\text{ mod }e\notin\{0,e/2\}$. Thus, that assumption leads to a contradiction and is therefore not valid. Instead, we employ approximation Eq. (S29) (detailed below) and take $\delta^{z}_{0}=-\delta^{z}_{3}=-d$ and $\delta^{z}_{1}=-\delta^{z}_{2}=0$; we assign the spin-up orbital associated with the bottom (top) surface of the bottom (top) layer the position $z=-d$ ($z=d$) and the spin-up orbital associated with the top (bottom) surface of the bottom (top) layer the position $z=0$ in the unit cell whose center is taken to be the point in the middle of the two layers (which are stacked on top of one another, see Fig. 1 of Ref. [30]). With this, we semi-analytically compute (S16), the results of which are in perfect agreement with Fig. 2 (a) of the main text. It is notable that in this case the calculation of $P^{z}$ is sensitive to the approximations made to the basis functions, while later in calculations of the topological magnetoelectric coefficient, this is no longer the case.
S1.3 Antiferromagnetic thin films
@@
$H_{\text{1L}}^{n=0}=\left(\begin{array}[]{cccc}m_{t}&\Delta_{\text{S}}\\ \Delta_{\text{S}}&m_{b}\\ \end{array}\right)=\frac{m_{t}+m_{b}}{2}\left(\begin{array}[]{cccc}1&0\\ 0&1\\ \end{array}\right)+\left(\begin{array}[]{cccc}(m_{t}-m_{b})/2&\Delta_{\text{S}}\\ \Delta_{\text{S}}&-(m_{t}-m_{b})/2\\ \end{array}\right),$ (S17)
which may be written as $H_{\text{1L}}=\bar{m}\tau_{0}+\tilde{m}\tau_{z}+\Delta_{\text{S}}\tau_{x}$ if we define $\bar{m}\equiv(m_{t}+m_{b})/2$ and $\tilde{m}\equiv(m_{t}-m_{b})/2$. Here $\tau_{i}$ denote the Pauli matrices. On the outside surfaces $m_{t/b}=\pm J_{\text{S}}$ whereas on the interior surfaces $m_{t/b}=\pm(1-\delta)J_{\text{S}}$ with $\delta\equiv J_{\text{D}}/J_{\text{S}}$. In the limit of $\delta\rightarrow 0$, the Hamiltonian in the $\mathcal{N}=0$ subspace have for a two layer thin film is:
$H_{\text{2L}}^{n=0}=\left(\begin{array}[]{cccc}J_{\text{S}}&\Delta_{\text{S}}&0&0\\ \Delta_{\text{S}}&J_{\text{S}}&0&0\\ 0&0&-J_{\text{S}}&\Delta_{\text{S}}\\ 0&0&\Delta_{\text{S}}&-J_{\text{S}}\\ \end{array}\right).$ (S18)
@@
In the case of $\delta\neq 0$, the layers in $N$-layer thin films may be grouped into the outer layers and interior layers. As an illustration, the following Hamiltonian describes a 4-layer thin film:
$H^{n=0}_{\text{4L}}=\left(\begin{array}[]{cccccccc}J_{\text{S}}&\Delta_{\text{S}}&0&0&0&0&0&0\\ \Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&0&0&0&0&0&0\\ 0&0&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}&0&0&0&0\\ 0&0&\Delta_{\text{S}}&-J_{\text{S}}+J_{\text{D}}&0&0&0&0\\ 0&0&0&0&J_{\text{S}}-J_{\text{D}}&\Delta_{\text{S}}&0&0\\ 0&0&0&0&\Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&0&0\\ 0&0&0&0&0&0&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}\\ 0&0&0&0&0&0&\Delta_{\text{S}}&-J_{\text{S}}\\ \end{array}\right).$ (S19)
Generalized to $N$-layer thin films, the energy eigenvalues and (non-normalized) eigenstates can be solved as:
@@
$E_{2N}=-(1-\delta/2)J_{\text{S}}-\sqrt{\Delta_{\text{S}}^{2}+(J_{\text{S}}\delta/2)^{2}};$ $\Psi_{2N}=\big{(}0,0,0,0,\cdots,0,0,\phi_{-},1\big{)}.$ (S20)
@@
As shown in the main text, the magnetoelectric coefficient may be written as:
$\alpha_{\text{me}}=\frac{e^{2}}{N(d_{s}+d_{v})hc}\sum_{j,E<E_{F}}|C_{E}(z_{j})|^{2}z_{j}.$ (S21)
@@
    For $J_{\text{S}}<\Delta_{\text{S}}$, $E_{2n}<0$ in Eq. (S1.3) and thus the contribution of interior layers is 0 since $|C_{E}(z_{j})|^{2}=C_{E}(-z_{j})|^{2}$.
    $\alpha_{\text{me}}|_{J_{\text{S}}<\Delta_{\text{S}}}=\frac{e^{2}}{Nhc}\frac{d_{s}}{d_{s}+d_{v}}\mathcal{A}_{-}^{2}(|\phi_{-}|^{2}-1),$ (S22)
@@
    Figure S1: (a) Thin film phase diagram for $J_{\text{S}}=1.5\Delta_{\text{S}}$ as discussed in the main text. (b)-(d) are the plots of the magnetoelectric coefficient versus the thickness of thin films for various $\delta$ and $\Delta_{\text{D}}$, with the values of $\delta$ labeled with black dashed lines in (a).
@@
      $\alpha_{\text{me}}^{\text{out}}=2\frac{e^{2}}{hc}\frac{N-1}{2N}+\frac{d_{s}}{2N(d_{s}+d_{v})}(\mathcal{A}_{+}^{2}+\mathcal{A}_{-}^{2}-\mathcal{A}_{+}^{2}\phi_{+}^{2}-\mathcal{A}_{-}^{2}\phi_{-}^{2}),$ (S23)
@@
      $\alpha_{\text{me}}|_{J_{\text{S}}>\Delta_{\text{S}},\delta<1-\Delta_{\text{S}}/J_{\text{S}}}=\alpha_{\text{me}}^{\text{out}}+\sum_{j=1}^{N/2-1}(-1)^{N/2+j}\alpha_{\text{me}}^{j}.$ (S24)
@@
      $\alpha_{\text{me}}|_{J_{\text{S}}>\Delta_{\text{S}},\delta\to 0}=\frac{e^{2}}{2hc}\sum_{j=1}^{N/2}(-1)^{N/2+j}\frac{2j-1}{N/2}.$ (S25)
@@
      $\alpha_{\text{me}}=\alpha_{\text{me}}^{\text{out}}=2\frac{e^{2}}{hc}\frac{N-1}{2N}+\frac{d_{s}}{2N(d_{s}+d_{v})}(\mathcal{A}_{+}^{2}+\mathcal{A}_{-}^{2}-\mathcal{A}_{+}^{2}\phi_{+}^{2}-\mathcal{A}_{-}^{2}\phi_{-}^{2}),$ (S26)
@@
For general values of $\delta$ and $\Delta_{\text{D}}$, the magnetoelectric coefficients calculated from numerical methods are shown in Fig. S1 (b)-(d), and the corresponding phase is shown in (a).
Appendix S2 Bulk calculations for three-dimensional MBX-type crystalline insulators
S2.1 A tight binding model
We employ a previously developed [15] square lattice regularization for the effective low-energy coupled Dirac cone model that was introduced (Eq. (2)) in the main text. The family of materials we consider are layered compounds (with stacking direction taken to be $\bm{z}$ and interlayer spacing $d$), and the tight-binding model we use was obtained by developing a 2D square lattice regularization of the $\bm{k}\cdot\bm{p}$ model of an isolated layer then coupling the states associated with the nearest surfaces of adjacent layers. In the non-magnetic case, the model is periodic under translations by $d\bm{z}$ (i.e., one layer per unit cell), while in the AFM case the model is periodic under translations by $2d\bm{z}$ (i.e., two layers per unit cell). The 3D square lattice regularized model (with Bravais lattice $\Gamma=\mathbb{Z}\times\mathbb{Z}\times 2d\mathbb{Z}$ and whose dual lattice is $\Gamma^{*}=2\pi\mathbb{Z}\times 2\pi\mathbb{Z}\times\frac{2\pi}{2d}\mathbb{Z}$) can be written in the general form
$\hat{H}^{(\text{3D})}_{\text{reg}}=\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\hat{c}^{\dagger}(\bm{k})\mathcal{H}^{(\text{3D})}_{\text{reg}}(\bm{k})\hat{c}(\bm{k}),$ (S27)
@@
$\hat{c}^{\dagger}(\bm{k})$ $\equiv\left(\hat{c}^{\dagger}_{(0,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(0,\downarrow),\bm{k}},\hat{c}^{\dagger}_{(1,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(1,\downarrow),\bm{k}},\hat{c}^{\dagger}_{(2,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(2,\downarrow),\bm{k}},\hat{c}^{\dagger}_{(3,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(3,\downarrow),\bm{k}}\right).$ (S28)
@@
$\frac{\partial}{\partial k^{a}}\bar{u}_{(\alpha,\sigma),\bm{k}}(\bm{r})=\sqrt{\frac{\Omega_{uc}}{(2\pi)^{3}}}\sum_{\bm{R}}\frac{\partial e^{-i\bm{k}\cdot(\bm{r}-\bm{R})}}{\partial k^{a}}W_{(\alpha,\sigma),\bm{R}}(\bm{r})\approx\sqrt{\frac{\Omega_{uc}}{(2\pi)^{3}}}\sum_{\bm{R}}\frac{\partial e^{-i\bm{k}\cdot\bm{\delta}_{\alpha}}}{\partial k^{a}}W_{(\alpha,\sigma),\bm{R}}(\bm{r})=-i\delta_{\alpha}^{a}\bar{u}_{(\alpha,\sigma),\bm{k}}(\bm{r}).$ (S29)
In the non-magnetic case, there is one layer per unit cell and if the layer thickness is neglected then one can choose $\bm{\delta}_{\alpha}=\bm{0}$, which is equivalent to the assumption of atomic-like WFs. However, in the AFM case there are two layers per unit cell, thus there is always more than one position within it about which WFs are localized. In particular, even if the WFs within each layer are taken to be localized about the same point, $\bm{\delta}_{0}=\bm{\delta}_{1}$ and $\bm{\delta}_{2}=\bm{\delta}_{3}$, we have $\bm{\delta}_{2}-\bm{\delta}_{0}=d\bm{z}$. In this approximation, the most natural choice is to take $\bm{\delta}_{2}=\frac{d}{2}\bm{z}$ and $\bm{\delta}_{0}=-\frac{d}{2}\bm{z}$.
@@
$\left(\begin{array}[]{cccccccc}\mu&iA(\text{s}k_{x}-i\text{s}k_{y})&\Delta_{\text{S}}(k_{x},k_{y})&0&0&0&e^{-2idk_{z}}\Delta_{\text{D}}&0\\ -iA(\text{s}k_{x}+i\text{s}k_{y})&-\mu&0&\Delta_{\text{S}}(k_{x},k_{y})&0&0&0&e^{-2idk_{z}}\Delta_{\text{D}}\\ \Delta_{\text{S}}(k_{x},k_{y})&0&\mu&-iA(\text{s}k_{x}-i\text{s}k_{y})&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}(k_{x},k_{y})&iA(\text{s}k_{x}+i\text{s}k_{y})&-\mu&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&-\mu&iA(\text{s}k_{x}-i\text{s}k_{y})&\Delta_{\text{S}}(k_{x},k_{y})&0\\ 0&0&0&\Delta_{\text{D}}&-iA(\text{s}k_{x}+i\text{s}k_{y})&\mu&0&\Delta_{\text{S}}(k_{x},k_{y})\\ e^{2idk_{z}}\Delta_{\text{D}}&0&0&0&\Delta_{\text{S}}(k_{x},k_{y})&0&-\mu&-iA(\text{s}k_{x}-i\text{s}k_{y})\\ 0&e^{2idk_{z}}\Delta_{\text{D}}&0&0&0&\Delta_{\text{S}}(k_{x},k_{y})&iA(\text{s}k_{x}+i\text{s}k_{y})&\mu\end{array}\right).$ (S38)
Here $\Delta_{\text{S}}(k_{x},k_{y})\equiv\Delta_{\text{S}}-2B(2-\text{c}k_{x}-\text{c}k_{y})$, where $\text{c}k_{x}\equiv\cos(k_{x})$, $\text{s}k_{x}\equiv\sin(k_{x})$, etc. This Hamiltonian is valid only when $\mu\neq 0$, since otherwise the material is non-magnetic and the crystalline unit cell is misidentified [15]. The eigenvalues of (S38) are
@@
$E_{3,4}(\bm{k})$ $=-E_{5,6}(\bm{k})=-\sqrt{\frac{1}{2}A^{2}(2-\text{c}2k_{x}-\text{c}2k_{y})+\Delta_{\text{S}}(k_{x},k_{y})^{2}+\Delta_{\text{D}}^{2}+\mu^{2}-2\sqrt{\Delta_{\text{S}}(k_{x},k_{y})^{2}(\Delta_{\text{D}}^{2}\cos^{2}(dk_{z})+\mu^{2})}}.$ (S39)
We will consider the ground state of the material to be that for which the energy bands are half-filled. We restrict our focus to band insulators, which implies $\Delta_{\text{S}}^{2}+\Delta_{\text{D}}^{2}+\mu^{2}-2\sqrt{\Delta_{\text{S}}^{2}(\Delta_{\text{D}}^{2}+\mu^{2})}>0$ (i.e., $|\Delta_{\text{S}}|-\sqrt{\Delta_{\text{D}}^{2}+\mu^{2}}\neq 0$). Moreover, the family of MBX materials of interest are such that the band gap is smallest at $\bm{k}=(0,0,0)$, which is always the case when $B/\Delta_{\text{S}}<0$. The double degeneracy of the energy bands at each $\bm{k}\in\text{BZ}$ follows from the combination of an inversion and a (fermionic) time-reversal symmetry (see Appendix C of Ref. [15]). A consequence is that the corresponding eigenvectors are highly non-unique. We wish to identify a set of orthogonal energy eigenvectors that are smooth and orthogonal over BZ such that they constitute a smooth global (and periodic) gauge choice of the bundle of occupied Bloch states over the Brillouin zone torus and of the total Bloch bundle more generally. One such choice, written in the basis (|$\big{(}\ket{\bar{\psi}_{(\alpha,\uparrow),\bm{k}}},\,\ket{\bar{\psi}_{(\alpha,\downarrow),\bm{k}}}\big{)}_{\alpha\in\{0,1,2,3\}}$, is

### v4

@@ A.2 Bulk polarization of the electronic ground state in the low-energy model
In the main text we demonstrated that, due to the symmetry of Eq. (18), the $\mathcal{N}>0$ Landau level subspaces do not contribute to the electronic polarization. Thus we restrict focus to the Hamiltonian in the $\mathcal{N}=0$ subspace, Eq. (21), which takes the form of a generalized SSH model. That is, when we consider an infinite number of layers (i.e. implement periodic boundary conditions) in Eq. (10), the result takes a tight-binding form [15] in the stacking ($\hat{\bm{z}}$) direction. And when this 3D effective model is projected into the $\mathcal{N}=0$ subspace, the result is a one-dimensional (1D) tight-binding model. The ground state polarization in our low-energy 3D model has one bulk component, namely $P^{z}$, which can be calculated using expressions derived in the modern theory of polarization [38, 39]. The $\mathcal{N}=0$ contribution to $P^{z}$ is proportional to the 1D bulk polarization of Eq. (21) at half-filling, namely
$P_{\text{SSH}}^{z}=e\sum_{n=1}^{2}\int_{\text{BZ}}\frac{dk_{z}}{2\pi}\xi^{z}_{nn}(k_{z}),$ (25)
@@
where $\xi^{z}_{nm}(k_{z})\equiv i\innerproduct{u_{n,k_{z}}}{\frac{\partial}{\partial k^{z}}u_{m,k_{z}}}$ are band components of the Berry connection induced by a choice of Bloch energy eigenvectors with cell-periodic parts |$\ket{u_{n,k_{z}}}$. Inversion symmetry of the model dictates that $P_{\text{SSH}}^{z}\text{ mod }e\in\{0,e/2\}$ [57]. A set of analytical eigenvectors of Eq. (21) can be found and the components differentiated with respect to $k_{z}$ in an effort to obtain $\xi^{z}_{nn}$.
@@
One issue remains: how do we take the $k_{z}$-derivative of the basis states with respect to which the Hamiltonian matrix (21) is written? Typically in phenomenological tight-binding models those basis states are taken to be such that the corresponding orbitals are highly localized in real-space and centered at the ionic positions. Often the derivatives of these basis states do not affect the results of calculations of topological quantities (e.g., polarization of the usual SSH model or the Chern number of the Haldane model) and one can get away with taking the basis states to be independent of $\bm{k}$. However, in the present case accounting for these derivatives becomes important. For if we assume the basis states to be independent of $k_{z}$ then we find, in general, that $P_{\text{SSH}}^{z}\text{ mod }e\notin\{0,e/2\}$. Thus, that assumption leads to a contradiction and is therefore not valid. Instead, we employ approximation Eq. (38) (detailed below) and take $\delta^{z}_{0}=-\delta^{z}_{3}=-d$ and $\delta^{z}_{1}=-\delta^{z}_{2}=0$; we assign the spin-up orbital associated with the bottom (top) surface of the bottom (top) layer the position $z=-d$ ($z=d$) and the spin-up orbital associated with the top (bottom) surface of the bottom (top) layer the position $z=0$ in the unit cell whose center is taken to be the point in the middle of the two layers (which are stacked on top of one another, see Fig. 1 of Ref. [31]). With this, we semianalytically compute Eq. (25), the results of which are in perfect agreement with Fig. 2 (a) of the main text. It is notable that in this case the calculation of $P^{z}$ is sensitive to the approximations made to the basis functions, while later in calculations of the topological magnetoelectric coefficient, this is no longer the case.
A.3 Antiferromagnetic thin films
@@
$H_{\text{1L}}^{n=0}=\left(\begin{array}[]{cccc}m_{t}&\Delta_{\text{S}}\\ \Delta_{\text{S}}&m_{b}\\ \end{array}\right)=\frac{m_{t}+m_{b}}{2}\left(\begin{array}[]{cccc}1&0\\ 0&1\\ \end{array}\right)+\left(\begin{array}[]{cccc}(m_{t}-m_{b})/2&\Delta_{\text{S}}\\ \Delta_{\text{S}}&-(m_{t}-m_{b})/2\\ \end{array}\right),$ (26)
which may be written as $H_{\text{1L}}=\bar{m}\tau_{0}+\tilde{m}\tau_{z}+\Delta_{\text{S}}\tau_{x}$ if we define $\bar{m}\equiv(m_{t}+m_{b})/2$ and $\tilde{m}\equiv(m_{t}-m_{b})/2$. Here $\tau_{i}$ denote the Pauli matrices. On the outside surfaces $m_{t/b}=\pm J_{\text{S}}$ whereas on the interior surfaces $m_{t/b}=\pm(1-\delta)J_{\text{S}}$ with $\delta\equiv J_{\text{D}}/J_{\text{S}}$. In the limit of $\delta\rightarrow 0$, the Hamiltonian in the $\mathcal{N}=0$ subspace have for a two layer thin film is
$H_{\text{2L}}^{n=0}=\left(\begin{array}[]{cccc}J_{\text{S}}&\Delta_{\text{S}}&0&0\\ \Delta_{\text{S}}&J_{\text{S}}&0&0\\ 0&0&-J_{\text{S}}&\Delta_{\text{S}}\\ 0&0&\Delta_{\text{S}}&-J_{\text{S}}\\ \end{array}\right).$ (27)
@@
In the case of $\delta\neq 0$, the layers in $N$-layer thin films may be grouped into the outer layers and interior layers. As an illustration, the following Hamiltonian describes a four-layer thin film:
$H^{n=0}_{\text{4L}}=\left(\begin{array}[]{cccccccc}J_{\text{S}}&\Delta_{\text{S}}&0&0&0&0&0&0\\ \Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&0&0&0&0&0&0\\ 0&0&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}&0&0&0&0\\ 0&0&\Delta_{\text{S}}&-J_{\text{S}}+J_{\text{D}}&0&0&0&0\\ 0&0&0&0&J_{\text{S}}-J_{\text{D}}&\Delta_{\text{S}}&0&0\\ 0&0&0&0&\Delta_{\text{S}}&J_{\text{S}}-J_{\text{D}}&0&0\\ 0&0&0&0&0&0&-J_{\text{S}}+J_{\text{D}}&\Delta_{\text{S}}\\ 0&0&0&0&0&0&\Delta_{\text{S}}&-J_{\text{S}}\\ \end{array}\right).$ (28)
Generalized to $N$-layer thin films, the energy eigenvalues and (nonnormalized) eigenstates can be solved as
@@
$E_{2N}=-(1-\delta/2)J_{\text{S}}-\sqrt{\Delta_{\text{S}}^{2}+(J_{\text{S}}\delta/2)^{2}};$ $\Psi_{2N}=\big{(}0,0,0,0,\cdots,0,0,\phi_{-},1\big{)}.$ (29)
@@
As shown in the main text, the magnetoelectric coefficient may be written as
$\alpha_{\text{me}}=\frac{e^{2}}{N(d_{s}+d_{v})hc}\sum_{j,E<E_{F}}|C_{E}(z_{j})|^{2}z_{j}.$ (30)
@@
    For $J_{\text{S}}<\Delta_{\text{S}}$, $E_{2n}<0$ in Eq. (A.3) and thus the contribution of interior layers is 0 since $|C_{E}(z_{j})|^{2}=C_{E}(-z_{j})|^{2}$.
    $\alpha_{\text{me}}|_{J_{\text{S}}<\Delta_{\text{S}}}=\frac{e^{2}}{Nhc}\frac{d_{s}}{d_{s}+d_{v}}\mathcal{A}_{-}^{2}(|\phi_{-}|^{2}-1),$ (31)
@@
      $\alpha_{\text{me}}^{\text{out}}=2\frac{e^{2}}{hc}\frac{N-1}{2N}+\frac{d_{s}}{2N(d_{s}+d_{v})}(\mathcal{A}_{+}^{2}+\mathcal{A}_{-}^{2}-\mathcal{A}_{+}^{2}\phi_{+}^{2}-\mathcal{A}_{-}^{2}\phi_{-}^{2}),$ (32)
@@
      $\alpha_{\text{me}}|_{J_{\text{S}}>\Delta_{\text{S}},\delta<1-\Delta_{\text{S}}/J_{\text{S}}}=\alpha_{\text{me}}^{\text{out}}+\sum_{j=1}^{N/2-1}(-1)^{N/2+j}\alpha_{\text{me}}^{j}.$ (33)
@@
      $\alpha_{\text{me}}|_{J_{\text{S}}>\Delta_{\text{S}},\delta\to 0}=\frac{e^{2}}{2hc}\sum_{j=1}^{N/2}(-1)^{N/2+j}\frac{2j-1}{N/2}.$ (34)
@@
      $\alpha_{\text{me}}=\alpha_{\text{me}}^{\text{out}}=2\frac{e^{2}}{hc}\frac{N-1}{2N}+\frac{d_{s}}{2N(d_{s}+d_{v})}(\mathcal{A}_{+}^{2}+\mathcal{A}_{-}^{2}-\mathcal{A}_{+}^{2}\phi_{+}^{2}-\mathcal{A}_{-}^{2}\phi_{-}^{2}),$ (35)
@@
Figure 4: (a) Thin film phase diagram for $J_{\text{S}}=1.5\Delta_{\text{S}}$ as discussed in the main text. Panels (b)-(d) are the plots of the magnetoelectric coefficient versus the thickness of thin films for various $\delta$ and $\Delta_{\text{D}}$, with the values of $\delta$ labeled with black dashed lines in panel (a).
For general values of $\delta$ and $\Delta_{\text{D}}$, the magnetoelectric coefficients calculated from numerical methods are shown in Figs. 4(b)-4(d), and the corresponding phase is shown in Fig. 4(a).
Appendix B Bulk calculations for three-dimensional MBX-type crystalline insulators
B.1 Tight-binding model
We employ a previously developed [15] square lattice regularization for the effective low-energy coupled Dirac cone model that was introduced (Eq. (2)) in the main text. The family of materials we consider are layered compounds (with stacking direction taken to be $\bm{z}$ and interlayer spacing $d$), and the tight-binding model we use was obtained by developing a 2D square lattice regularization of the $\bm{k}\cdot\bm{p}$ model of an isolated layer then coupling the states associated with the nearest surfaces of adjacent layers. In the nonmagnetic case, the model is periodic under translations by $d\bm{z}$ (i.e., one layer per unit cell), while in the AFM case the model is periodic under translations by $2d\bm{z}$ (i.e., two layers per unit cell). The 3D square lattice regularized model (with Bravais lattice $\Gamma=\mathbb{Z}\times\mathbb{Z}\times 2d\mathbb{Z}$ and whose dual lattice is $\Gamma^{*}=2\pi\mathbb{Z}\times 2\pi\mathbb{Z}\times\frac{2\pi}{2d}\mathbb{Z}$) can be written in the general form
$\hat{H}^{(\text{3D})}_{\text{reg}}=\int_{\text{BZ}}\frac{d^{3}k}{(2\pi)^{3}}\hat{c}^{\dagger}(\bm{k})\mathcal{H}^{(\text{3D})}_{\text{reg}}(\bm{k})\hat{c}(\bm{k}),$ (36)
@@ B.1 Tight-binding model
$\hat{c}^{\dagger}(\bm{k})$ $\equiv\left(\hat{c}^{\dagger}_{(0,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(0,\downarrow),\bm{k}},\hat{c}^{\dagger}_{(1,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(1,\downarrow),\bm{k}},\hat{c}^{\dagger}_{(2,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(2,\downarrow),\bm{k}},\hat{c}^{\dagger}_{(3,\uparrow),\bm{k}},\hat{c}^{\dagger}_{(3,\downarrow),\bm{k}}\right).$ (37)
@@
$\frac{\partial}{\partial k^{a}}\bar{u}_{(\alpha,\sigma),\bm{k}}(\bm{r})=\sqrt{\frac{\Omega_{uc}}{(2\pi)^{3}}}\sum_{\bm{R}}\frac{\partial e^{-i\bm{k}\cdot(\bm{r}-\bm{R})}}{\partial k^{a}}W_{(\alpha,\sigma),\bm{R}}(\bm{r})\approx\sqrt{\frac{\Omega_{uc}}{(2\pi)^{3}}}\sum_{\bm{R}}\frac{\partial e^{-i\bm{k}\cdot\bm{\delta}_{\alpha}}}{\partial k^{a}}W_{(\alpha,\sigma),\bm{R}}(\bm{r})=-i\delta_{\alpha}^{a}\bar{u}_{(\alpha,\sigma),\bm{k}}(\bm{r}).$ (38)
In the nonmagnetic case, there is one layer per unit cell and if the layer thickness is neglected then one can choose $\bm{\delta}_{\alpha}=\bm{0}$, which is equivalent to the assumption of atomic-like WFs. However, in the AFM case there are two layers per unit cell, thus there is always more than one position within it about which WFs are localized. In particular, even if the WFs within each layer are taken to be localized about the same point, $\bm{\delta}_{0}=\bm{\delta}_{1}$ and $\bm{\delta}_{2}=\bm{\delta}_{3}$, we have $\bm{\delta}_{2}-\bm{\delta}_{0}=d\bm{z}$. In this approximation, the most natural choice is to take $\bm{\delta}_{2}=\frac{d}{2}\bm{z}$ and $\bm{\delta}_{0}=-\frac{d}{2}\bm{z}$.
@@
$\left(\begin{array}[]{cccccccc}\mu&iA(\text{s}k_{x}-i\text{s}k_{y})&\Delta_{\text{S}}(k_{x},k_{y})&0&0&0&e^{-2idk_{z}}\Delta_{\text{D}}&0\\ -iA(\text{s}k_{x}+i\text{s}k_{y})&-\mu&0&\Delta_{\text{S}}(k_{x},k_{y})&0&0&0&e^{-2idk_{z}}\Delta_{\text{D}}\\ \Delta_{\text{S}}(k_{x},k_{y})&0&\mu&-iA(\text{s}k_{x}-i\text{s}k_{y})&\Delta_{\text{D}}&0&0&0\\ 0&\Delta_{\text{S}}(k_{x},k_{y})&iA(\text{s}k_{x}+i\text{s}k_{y})&-\mu&0&\Delta_{\text{D}}&0&0\\ 0&0&\Delta_{\text{D}}&0&-\mu&iA(\text{s}k_{x}-i\text{s}k_{y})&\Delta_{\text{S}}(k_{x},k_{y})&0\\ 0&0&0&\Delta_{\text{D}}&-iA(\text{s}k_{x}+i\text{s}k_{y})&\mu&0&\Delta_{\text{S}}(k_{x},k_{y})\\ e^{2idk_{z}}\Delta_{\text{D}}&0&0&0&\Delta_{\text{S}}(k_{x},k_{y})&0&-\mu&-iA(\text{s}k_{x}-i\text{s}k_{y})\\ 0&e^{2idk_{z}}\Delta_{\text{D}}&0&0&0&\Delta_{\text{S}}(k_{x},k_{y})&iA(\text{s}k_{x}+i\text{s}k_{y})&\mu\end{array}\right).$ (47)
Here $\Delta_{\text{S}}(k_{x},k_{y})\equiv\Delta_{\text{S}}-2B(2-\text{c}k_{x}-\text{c}k_{y})$, where $\text{c}k_{x}\equiv\cos(k_{x})$, $\text{s}k_{x}\equiv\sin(k_{x})$, etc. This Hamiltonian is valid only when $\mu\neq 0$, since otherwise the material is nonmagnetic and the crystalline unit cell is misidentified [15]. The eigenvalues of Eq. (47) are
@@
$E_{3,4}(\bm{k})$ $=-E_{5,6}(\bm{k})=-\sqrt{\frac{1}{2}A^{2}(2-\text{c}2k_{x}-\text{c}2k_{y})+\Delta_{\text{S}}(k_{x},k_{y})^{2}+\Delta_{\text{D}}^{2}+\mu^{2}-2\sqrt{\Delta_{\text{S}}(k_{x},k_{y})^{2}(\Delta_{\text{D}}^{2}\cos^{2}(dk_{z})+\mu^{2})}}.$ (48)
We will consider the ground state of the material to be that for which the energy bands are half-filled. We restrict our focus to band insulators, which implies $\Delta_{\text{S}}^{2}+\Delta_{\text{D}}^{2}+\mu^{2}-2\sqrt{\Delta_{\text{S}}^{2}(\Delta_{\text{D}}^{2}+\mu^{2})}>0$ (i.e., $|\Delta_{\text{S}}|-\sqrt{\Delta_{\text{D}}^{2}+\mu^{2}}\neq 0$). Moreover, the family of MBX materials of interest are such that the band gap is smallest at $\bm{k}=(0,0,0)$, which is always the case when $B/\Delta_{\text{S}}<0$. The double degeneracy of the energy bands at each $\bm{k}\in\text{BZ}$ follows from the combination of an inversion and a (fermionic) time-reversal symmetry (see Appendix C of Ref. [15]). A consequence is that the corresponding eigenvectors are highly nonunique. We wish to identify a set of orthogonal energy eigenvectors that are smooth and orthogonal over BZ such that they constitute a smooth global (and periodic) gauge choice of the bundle of occupied Bloch states over the Brillouin zone torus and of the total Bloch bundle more generally. One such choice, written in the basis (|$\big{(}\ket{\bar{\psi}_{(\alpha,\uparrow),\bm{k}}},\,\ket{\bar{\psi}_{(\alpha,\downarrow),\bm{k}}}\big{)}_{\alpha\in\{0,1,2,3\}}$, is